# Phase transitions at high and low densities for a rotating QCD matter from holography

### v1

@@ 8 Estimation for the critical point of the QCD phase diagram
The IR parameter $\sqrt{c}$ can be fixed in the soft wall model using QCD phenomenology. The fit of the masses of lightest $\rho$-mesons leads to $\sqrt{c}=338\,\text{MeV}$ [59]. The $\eta$ parameter, which appears in the relation between the BH charge parameter and the quark chemical potential, see (2.16), affects the HP transitions in a non-trivial way [39]. For simplicity, we have assumed $\eta=1$. To recover its dependence, one must define the physical quark density
@@
The baryon chemical potentials are given by $\mu_{B}=3\mu$, see [39], so that the holographic prediction for the baryon density at CP is
$\mu_{CPB}\approx 363.554\,\text{MeV}\;.$ (8.6)
@@
$\mu_{0B}\approx 937.425\,\text{MeV}\;,$ (8.8)
such that $\mu_{CPB}/\mu_{0B}\approx 38.8\%$, which shows that the smooth transitions must occur before the first half of the QCD phase diagram.
@@ 9 Conclusions
In this work, we have studied the phase structure of rotating QCD matter within Andreev’s soft-wall holographic model, considering a charged BH with nonzero angular momentum in AdS5 spacetime. The Hawking-Page transitions were analyzed to describe the deconfinement of hadronic matter into a QGP at finite density and rotation. Our results reveal a rich interplay between chemical potential, temperature, and angular velocity. For quark chemical potentials above the maximum $\bar{\mu}_{EF}$, the phase transitions are predominantly of first order. The critical temperature decreases with increasing chemical potential, and the transitions remain sharp, reflecting conventional confinement-deconfinement behavior. In the non-relativistic limit, corresponding to small rotational velocities ($\omega l\lesssim 0.16$), the transitions are well-described by the critical temperature $T_{c}(\omega l\approx 0)$, and the coexistence of phases is negligible, even at intermediate densities. The second phase transition that may occur at extremely high temperatures corresponds to highly unstable hadronic states, and it is disregarded, as the free energy analysis loses validity in this regime due to the vanishing QCD coupling constant.
@@
On the other hand, for chemical potentials below $\text{max}(\bar{\mu}_{EF})\approx 0.414$, relativistic rotations play a significant role in shaping the phase diagram. At high temperatures and low densities, where kinetic energy is abundant, the QGP and hadronic matter coexist over a wide range of angular momenta. In this region, transitions occur smoothly rather than abruptly, leading to crossover-like behavior instead of first-order transitions. The coexistence region is influenced by the negative QCD coupling constant, and the total Gibbs free energy must be considered as an infinite sum over all coexisting states. These smooth transitions occur predominantly in intermediate low-density regions, while at lower temperatures, hadronic matter remains stable, and the transitions are controlled by the non-rotating critical temperature. Our analysis demonstrates that the relativistic effect of rotation is most pronounced for transitions labeled $E\rightarrow F$, occurring at high temperatures and low densities. At lower rotational velocities or in high-density, low-temperature regimes, the coexistence of phases is negligible, and first-order transitions dominate. Overall, the exact soft-wall model provides a robust framework to capture both conventional first-order transitions at high densities and rotationally induced smooth crossover transitions at low densities, highlighting the significant impact of angular momentum on the QCD phase diagram.
@@
In Fig. 11, we plot the QCD phase diagram including all results from the analysis of phase transitions at low and high densities of strongly interacting matter in Andreev’s exact soft-wall AdS/QCD model, accounting for relativistic rotations. Between the first-order Herzog-type transition at zero density and those occurring in the high-density regime, there is an intermediate region in which hadronic matter and plasma can coexist with different angular momenta. This intermediate region is characterized by crossover transitions, since it is described by $E\rightarrow F$ transitions, see Sec. 6, whose states vary smoothly in the phase diagram, rather than by jumping from one phase to another. This behavior appears in the low-density regime for rotating matter, as shown in Figs. 5 and 6, in which one can observe the existence of an interval where the HP transition no longer occurs, and the system is always in the confined phase. The critical chemical potential that defines this interval in which hadrons are always stable depends on the rotational velocity, see Fig. 7. For this reason, the holographic description indicates the existence of a mixed phase that can occur at low densities and high temperatures, as inferred from the analysis in Fig. 8. The subdivisions shown in this figure could be made infinitely, corresponding to smooth point-to-point transitions (crossovers), according to the phases mixing defined by the angular momentum of the particles.
@@
The interval in which the crossover transitions occur is defined by Eq. (8.1), as we can conclude from the analysis of Fig. 7. There is a maximum value of the chemical potential for which this type of coexistence between phases can occur for states with distinct angular momentum. This defines the critical point between smooth transitions at low densities and first-order transitions at high densities and lower temperatures. In terms of baryonic density, the prediction of Andreev’s holographic model is given by
$(\mu_{CPB},\,T_{CP})=(363.554,\,58.507)\,\text{MeV}\;,$ (9.1)
@@
which was obtained through numerical and phenomenological analysis described in Sec. 8. This CP is indicated in Fig. 11. For $\mu\geq\mu_{CPB}$, the system is described by first-order transitions, dominated by the $T_{c}(\omega l\approx 0)$ curve, for states with rotational velocities much lower than the speed of light, as discussed in Sec. 7. The distinction between the types of phase transitions at high and low densities is attributed to the effect of relativistic rotations, which can only occur for $\omega l\gtrsim 0.16$, in the low-density regime ($\mu\leq\mu_{CPB})$, see Fig. 7. This estimate indicates that this effect should be considerable in states at very high temperatures, as shown by the yellow-hatched region in Fig. 11. In the non-relativistic limit, the $E\to F$ transitions cease to occur and are automatically described by first-order transitions, in which hadron matter is more stable at lower temperatures, see Fig. 10. Meanwhile, at the limit of extremely high temperatures, the analysis of Gibbs free energies loses its power to describe the stability of states, since for $g^{2}\rightarrow 0$, both action densities (of the thermal and BH AdS geometries) diverge to infinity. In this case, the transition, which occurs smoothly as the temperature increases, should be governed solely by the negative QCD $\beta$-function, from a mixed state – QGP and hadrons with different angular momentum – to a QGP at extremely high temperatures.
@@
Building on the comprehensive analysis presented here of rotational effects on the QCD phase diagram within an exact soft-wall holographic model, several concrete perspectives emerge to further refine the understanding of rotating strongly interacting matter. A natural next step is to incorporate the full backreaction of the rotating charged geometry in AdS, enabling an even more precise determination of the interface between crossover behavior and first-order transitions. Incorporating subleading dilaton and higher-derivative corrections could test the stability of the coexistence window by revealing whether the smooth transition region persists once the AdS bulk dynamics is modified beyond the leading soft-wall approximation, or whether it collapses into a sharper first-order structure when higher-order contributions to the action and metric response are taken into account. Moreover, calculating two-point functions and quasinormal spectra for AdS bulk perturbations dual to both the hadronic phase and the QGP plasma states with different boundary angular momenta would clarify whether the mixed configurations identified here manifest distinct dynamical signatures. The specific heat, which quantifies how the system’s energy responds to temperature changes, and the baryon number susceptibility, which measures how baryon density responds to chemical potential, both serve as probes of the phase structure. Peaks or divergences in these quantities signal the approach to a critical point and can be used to independently locate the transition between crossover and first-order behavior. Together, these steps would refine the predictive power of the soft-wall model in the rotational regime of the QGP.

### v2

@@ 7 Estimation for the critical point of the QCD phase diagram
The IR parameter $\sqrt{c}$ can be fixed in the soft wall model using QCD phenomenology. The fit of the masses of lightest $\rho$-mesons leads to $\sqrt{c}=338\,\text{MeV}$ [^2] This phenomenological estimate of $\sqrt{c}$ was obtained using a zero-density holographic soft wall model with $A_{z}=0$, see [58]. We do not rule out a possible refinement of this value, assuming the effect of rotation (where $A_{z}\neq 0$) and of the quark chemical potential. [58]. The $\eta$ parameter, which appears in the relation between the BH charge parameter and the quark chemical potential, see (2.16), affects the HP transitions in a non-trivial way [38]. For simplicity, we have assumed $\eta=1$. To recover its dependence, one must define the physical quark density
@@
The baryon chemical potentials are given by $\mu_{B}=3\mu$, see [38], so that the holographic prediction for the baryon density at CP for $N_{f}=6$ is
$\mu_{CPB}(N_{f}=6)\approx 363.554\,\text{MeV}\;.$ (7.6)
This estimation is strongly affected by the number of flavors, since $\eta$ is inversely proportional to $\sqrt{N_{f}}$, see Eq. (7.4). For $N_{f}=3$, one finds
$\mu_{CPB}(N_{f}=3)\approx 514.143\,\text{MeV}\;.$ (7.7)
The chemical potential measured by an observer at a given frame tends to decrease as the rotational velocity increases. The estimations above are valid for an observer in a non-rotating frame. For the rotating one, we must use Eq. (2.18) with $\omega l\approx 0.3$ at CP, which yields
$\mu_{CPB}^{\prime}(N_{f}=6)$ $\approx$ $346.808\,\text{MeV}\;,$ (7.8)
$\mu_{CPB}^{\prime}(N_{f}=3)$ $\approx$ $460.461\,\text{MeV}\;.$ (7.9)
@@
$\mu_{0B}(N_{f}=6)$ $\approx$ $937.425\,\text{MeV}\;,$ (7.11)
$\mu_{0B}(N_{f}=3)$ $\approx$ $1325.100\,\text{MeV}\;,$ (7.12)
such that $\mu_{CPB}/\mu_{0B}\approx 38.8\%$, which shows that the smooth transitions must occur before the first half of the QCD phase diagram. The values of (7.11) and (7.12) are also valid for the rotating frame; this transition is dominated by the non-rotating $T_{c}$ curve with $\omega l\approx 0$. On the other hand, the results (7.8) and (7.9) are sensitive even to small changes in the rotational velocity, which were estimated numerically at the critical point, with an accuracy of $\omega l=0.30\pm\delta\omega$, with $\delta\omega=0.01$.
8 Comparison with lattice QCD, effective models and holography
The phase structure obtained in this work can be compared with results from lattice QCD, effective models, and previous holographic approaches, including rotation. Lattice simulations at finite angular velocity indicate that rotation tends to reduce the critical temperature of deconfinement, with the shift depending on the magnitude of the vorticity and the system size [81, 82, 83]. This qualitative behavior is consistent with Eq. (2.6), where the factor $\sqrt{1-\omega^{2}l^{2}}$ suppresses the temperature, leading to a decrease of $T_{c}$ with increasing rotation. Our results reproduce this tendency both at zero and finite density, and extend it by showing that, in the relativistic regime, rotation also modifies the order of the transition through the emergence of phase coexistence regions associated with $E\rightarrow F$ transitions.
In effective approaches such as NJL and PNJL models, rotation is known to affect chiral symmetry restoration and deconfinement, typically lowering the critical temperature and shifting the critical endpoint toward smaller chemical potentials [84, 85, 86]. The mechanism is usually attributed to rotational suppression of condensates and modification of the fermionic spectrum. Our holographic results are compatible with this trend: the decrease of $T_{c}$ with $\omega l$ and the existence of a finite critical chemical potential separating crossover and first-order regions are qualitatively similar. However, in contrast to NJL/PNJL models, the present approach provides a geometric interpretation in terms of competing bulk solutions. Naturally, it incorporates strong-coupling effects, thereby identifying a mixed phase driven by angular-momentum distributions rather than solely by chiral dynamics.
The predicted critical point can also be compared quantitatively with other approaches. In the present model, we obtain $(\mu_{B},T)_{CP}\approx(363\,\text{MeV}$$514\,\text{MeV},58\,\text{MeV})$. Lattice extrapolations typically suggest a critical point, if present, in the range $\mu_{B}\sim 300$ - $600\,\text{MeV}$ and $T\sim 100$ - $160\,\text{MeV}$, although large uncertainties remain [87, 88]. NJL and PNJL models often predict $T_{CP}$ between $50$ and $120\,\text{MeV}$ and $\mu_{B}$ between $300$ and $800\,\text{MeV}$, depending on parameter choices [85, 86]. Compared to these results, our prediction lies in the lower-temperature region but within the expected range of baryon chemical potentials. The relatively low value of $T_{CP}$ can be attributed to the soft-wall scale and to the strong suppression induced by relativistic rotation, which effectively shifts the crossover region toward smaller temperatures.
@@ 9 Conclusions
In this work, we have studied the phase structure of rotating QCD matter within Andreev’s soft-wall holographic model, considering a charged BH with nonzero angular momentum in AdS5 spacetime. The Hawking-Page transitions were analyzed to describe the deconfinement of hadronic matter into a QGP at finite density and rotation. Our results reveal a rich interplay between chemical potential, temperature, and angular velocity. For quark chemical potentials above the maximum $\bar{\mu}_{EF}$, the phase transitions are predominantly of first order. The critical temperature decreases with increasing chemical potential, and the transitions remain sharp, reflecting conventional confinement-deconfinement behavior. In the non-relativistic limit, corresponding to small rotational velocities ($\omega l\lesssim 0.16$), the transitions are well-described by the critical temperature $T_{c}(\omega l\approx 0)$, and the coexistence of phases is negligible, even at intermediate densities. The second phase transition that may occur at extremely high temperatures corresponds to highly unstable hadronic states and is disregarded, due to the vanishing QCD coupling constant and their thermodynamic properties.
@@
On the other hand, for chemical potentials below $\text{max}(\bar{\mu}_{EF})\approx 0.414$, relativistic rotations play a significant role in shaping the phase diagram. Assuming an ensemble with $N_{p}$ particles, being $N_{p}$ large, at high temperatures and low densities, where kinetic energy is abundant, the QGP and hadronic matter coexist over a wide range of angular momenta. In this region, transitions occur smoothly rather than abruptly, resulting in crossover-like behavior rather than first-order transitions. This interpretation can be explained from the analysis of the thermodynamic properties of the system (Gibbs free energy, entropy, specific heat) for each phase, which was carried out in Section 6, using holographic renormalization. The negative QCD coupling constant influences the coexistence region, and the total Gibbs free energy must be considered as an infinite sum over all coexisting states. These smooth transitions occur predominantly in intermediate-low-density regions, whereas at lower temperatures, hadronic matter remains stable and the transitions are governed by the non-rotating critical temperature. Our analysis demonstrates that the relativistic effect of rotation is most pronounced for transitions labeled $E\rightarrow F$, occurring at high temperatures and low densities. At lower rotational velocities or in high-density, low-temperature regimes, the coexistence of phases is negligible, and first-order transitions dominate. Overall, the exact soft-wall model provides a robust framework to capture both conventional first-order transitions at high densities and rotationally induced smooth crossover transitions at low densities, highlighting the significant impact of angular momentum on the QCD phase diagram.
@@
In Fig. 19, we plot the QCD phase diagram including all results from the analysis of phase transitions at low and high densities of strongly interacting matter in Andreev’s exact soft-wall AdS/QCD model, accounting for relativistic rotations. Between the first-order Herzog-type transition at zero density and those occurring in the high-density regime, there is an intermediate region in which hadronic matter and plasma can coexist with different angular momenta. Crossover transitions characterize this intermediate region, since it is described by $E\rightarrow F$ transitions, see Sec. 6, whose states vary smoothly in the phase diagram, rather than by jumping from one phase to another. This behavior is observed in the low-density regime of rotating matter, as shown in Figs. 5 and 6, in which one can observe the existence of an interval where the HP transition no longer occurs, and the system is always in the confined phase. The critical chemical potential that defines this interval in which hadrons are always stable depends on the rotational velocity, see Fig. 7. For this reason, the holographic description indicates the existence of a mixed phase that can occur at low densities and high temperatures, as inferred from the analysis in Fig. 8. The subdivisions shown in this figure could be made infinitely, corresponding to smooth point-to-point transitions (crossovers), according to the phases mixing defined by the angular momentum of the particles.
@@
The interval over which crossover transitions occur is defined by Eq. (7.1), as can be inferred from the analysis of Fig. 7. There is a maximum value of the chemical potential for which this type of coexistence between phases can occur for states with distinct angular momentum. This defines the critical point between smooth transitions at low densities and first-order transitions at high densities and lower temperatures. In terms of baryonic density, the prediction of Andreev’s holographic model in the non-rotating frame, for a system with $N_{f}=6$ flavors, is given by Eqs. (7.6, 7.9). These values were obtained through numerical and phenomenological analysis described in Sec. 7. The CP for $N_{f}=6$ is indicated in Fig. 17. For $\mu\geq\mu_{CPB}$, the system is described by first-order transitions, dominated by the $T_{c}(\omega l\approx 0)$ curve, for states with rotational velocities much lower than the speed of light, as discussed in Sec. 6.2. The distinction between the types of phase transitions at high and low densities is attributed to the effect of relativistic rotations, which can only occur for $\omega l\gtrsim 0.16$, in the low-density regime ($\mu\leq\mu_{CPB})$, see Fig. 7. This estimate indicates that this effect should be considerable in states at very high temperatures, as shown by the yellow-hatched region in Fig. 19. In the non-relativistic limit, the $E\to F$ transitions cease to occur and are automatically described by first-order transitions, in which hadron matter is more stable at lower temperatures, see Fig. 18. Meanwhile, at the limit of extremely high temperatures, it is expected that the hadronic matter will become unstable, due to the negative $\beta$-function of QCD. This instability is consistently described by the present holographic model, since the specific heat of hadronic states (dual to the thermal AdS geometry) becomes negative as the temperature increases, see Fig. 13, indicating a marked loss of energy of these states. In this case, the transition will occur smoothly, from a mixed state – QGP and hadrons with different angular momentum – to a phase dominated by QGP states at extremely high temperatures. On the other hand, at the low-temperature limit, hadronic matter is dominant. These transitions, from stable hadronic matter at lower temperatures to a mixed phase, and then from the mixed phase to a phase energetically dominated by QGP states, are consistently described by the total energy of the system in the grand-canonical ensemble, see Fig. 16. Such transitions are indeed smooth, as demonstrated by Figures 11 and 12, in which the entropy and specific heat curves for the ensemble do not indicate any discontinuities.
@@
Building on the comprehensive analysis presented here of rotational effects on the QCD phase diagram within an exact soft-wall holographic model, several concrete perspectives emerge to further refine the understanding of rotating strongly interacting matter. A natural next step is to incorporate the full backreaction of the rotating charged geometry in AdS, enabling an even more precise determination of the interface between crossover behavior and first-order transitions. Incorporating subleading dilaton and higher-derivative corrections could test the stability of the coexistence window by revealing whether the smooth transition region persists once the AdS bulk dynamics is modified beyond the leading soft-wall approximation, or whether it collapses into a sharper first-order structure when higher-order contributions to the action and metric response are taken into account. Moreover, calculating two-point functions and quasinormal spectra for AdS bulk perturbations dual to both the hadronic phase and the QGP plasma states with different boundary angular momenta would clarify whether the mixed configurations identified here manifest distinct dynamical signatures. The specific heat, which quantifies how the system’s energy responds to temperature changes, and the baryon number susceptibility, which measures how baryon density responds to chemical potential, both serve as probes of the phase structure. Peaks or divergences in these quantities signal the approach to a critical point and can be used to locate the transition between crossover and first-order behavior independently. Together, these steps would refine the predictive power of the soft-wall model in the QGP rotational regime.
@@ Appendix A Auxiliary tables with numerical results
The values of the variables in the tables are in their dimensionless versions. To obtain these values in MeV units, simply apply equations (2.23) and (3.1), with $\sqrt{c}=338\,\text{MeV}$.